OpenStudy (anonymous):

Factor: 4x2 + 8x + 3 A. (2x + 1)(2x + 3) B. (4x - 1)(x + 2) C. (4x + 3)(x + 1) D. prime

OpenStudy (arnavguddu):

4x2 + 8x + 3 = 4x2 + 6x + 2x + 3 = 2x(2x+3) + 1(2x+3) = (2x+1)(2x+3) Ans-A

OpenStudy (apoorvk):

A